OpenStudy (exoexoexo):

2 planes start at 11AM from two airports 1800 miles apart, and fly toward each other at average rates of 200 and 250 miles per hour. the time at which they pass each other will be?

OpenStudy (hba):

Lets say after 2 hours

OpenStudy (exoexoexo):

plane a 400 miles plane b 500 miles

OpenStudy (hba):

1 PM

OpenStudy (exoexoexo):

a 600 miles b 750 miles

OpenStudy (kropot72):

The closing speed of the two planes = 200 + 250 = 450 mph The combined distance traveled by the planes is 1800 miles before they pass each other is 1800 miles. Therefore the elapsed time before the planes pass is given by 1800/450 hours. Now after calculating the elapsed time you need to add it to 11 AM.

OpenStudy (kropot72):

@EXOEXOEXO Have you calculated 1800/450 yet?

OpenStudy (kropot72):

@EXOEXOEXO Are you there?

OpenStudy (exoexoexo):

yes

OpenStudy (exoexoexo):

sorry

OpenStudy (exoexoexo):

4 hours

OpenStudy (kropot72):

Have you calculated 1800/450 yet? Please post your result.

OpenStudy (exoexoexo):

1800/450=4 4 hours after 11am is 3pm

OpenStudy (kropot72):

You are correct!

OpenStudy (exoexoexo):

thank you!

OpenStudy (kropot72):

you're welcome :)